# Mills’ constant is irrational

### v1

@@ 3. Lemmas
Take any $k\in\mathbb{N}$. By the definition of $\xi$, we see that $p_{k}\leq\xi^{C_{k}}<p_{k}+1$, which implies that $p_{k}^{c_{k+1}}\leq\xi^{C_{k+1}}<(p_{k}+1)^{c_{k+1}}$. We recall that $c_{k+1}$ is a positive integer greater than or equal to $2$, and hence $p_{k}^{c_{k+1}}\leq\lfloor\xi^{C_{k+1}}\rfloor<(p_{k}+1)^{c_{k+1}}$. It is not true that $p_{k+1}=(p_{k}+1)^{c_{k+1}}-1$ since $p_{k+1}$ is a prime number. Therefore, we obtain (3.1). ∎
@@
There exists $k_{0}=k_{0}(\xi)>0$ such that for all $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$, we have
@@
To prove Lemma 3.2, we apply the following result given by Matomäki (see [11, Lemma 9] and [10]).
@@
There exist positive constants $d_{1}<1$ and $D$ such that, for every sufficiently large $x$ and every $\gamma\in[1/2,1]$, the interval $[x,2x]$ contains at most $Dx^{2/3-\gamma}$ disjoint intervals $[n,n+n^{\gamma}]$ for which
@@
By applying this theorem, Matomäki [11] constructed an element in $\mathcal{W}(2)$. It is truly astonishing because Mills’ construction tells us that we need the existence of a prime number in each interval $[n,n+n^{1/2}]$ to concrete an element in $\mathcal{W}(2)$. This problem is strongly connected with Legendre’s conjecture which asserts that for every $n$ there exists a prime number between $n^{2}$ and $(n+1)^{2}$. It is unsolved and believed to be extremely difficult. However, Theorem 3.3 with $\gamma=1/2$ states that there exists a prime number $p\in[n,n+n^{1/2}]$ except for very few intervals $[n,n+n^{1/2}]$. Therefore, Matomäki discovered that we can construct a PRC if the exceptional intervals are very few. We apply Theorem 3.3 to prove the following lemma.
@@
Let $c$ be a real number in the interval $[2,B]$, and $\epsilon$ be a sufficiently small positive real number. Let $X_{0}$ be a sufficiently large positive parameter depending only on $B$ and $\epsilon$. Let $X$ and $\eta$ be real numbers satisfying $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$. Suppose that
(3.3) #$\#([X,X+X^{\eta}]\cap\mathcal{P})\geq\frac{d_{2}X^{\eta}}{\log X}$
for some constant $d_{2}>0$. Then there exists $q\in[X,X+X^{\eta}]\cap\mathcal{P}$ such that
@@
Let $R=[X,X+X^{\eta}]\cap\mathcal{P}$. Each interval $[q^{c},q^{c}+q^{c-1}]$ $(q\in R)$ is completely contained in $[X^{c},2X^{c}]$. Indeed, we have
@@
$\leq X^{c}(1+X^{\eta-1})^{B}(1+X^{-1})\leq 2X^{c}$
since $\eta\leq 1-\epsilon$ and $X$ is sufficiently large. Setting $\gamma=(c-1)/c$, we have $\gamma\in[1/2,1)$. By $\eta\geq 1/2$ and $c\geq 2$, the following inequalities are true:
(3.5) $\frac{d_{2}X^{\eta}}{\log X}-DX^{c(2/3-\gamma)}$ $=\frac{d_{2}X^{\eta}}{\log X}-DX^{1-c/3}=\frac{X^{\eta}}{\log X}\left(d_{2}-DX^{1-\eta-c/3}\log X\right)$
$\geq\frac{{X^{\eta}}}{\log X}\left(d_{2}-DX^{-1/6}\log X\right)>0,$
@@
We now take $k_{0}$ as a sufficiently large positive parameter. By the assumption, we find an integer $k$ in $\mathcal{I}_{b}\cap[k_{0},\infty)$ satisfying (3.6). By Lemma 3.1, it is false that $p_{k}^{c_{k+1}}>p_{k+1}$, and hence we have
@@
#$\#([q_{k}^{c_{k+1}},q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}]\cap\mathcal{P})\geq\frac{d_{0}q_{k}^{\theta c_{k+1}}}{\log q_{k}^{c_{k+1}}}.$
By Lemma 3.4 with $X\coloneqq q_{k}^{c_{k+1}}$, $\eta\coloneqq\theta$, and $c\coloneqq c_{k+2}$, there exists $q_{k+1}\in\mathcal{P}$ satisfying
(3.8) $q_{k}^{c_{k+1}}\leq q_{k+1}\leq q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}(\leq q_{k}^{c_{k+1}}+q_{k}^{c_{k+1}-1})$
(3.9) #$\#([q_{k+1}^{c_{k+2}},q_{k+1}^{c_{k+2}}+q_{k+1}^{c_{k+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c_{k+2}-1}}{\log q^{c_{k+2}}}.$
Let us construct a sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying the following properties: for every $m\geq k$, we have
@@
(3.11) #$\#([q_{m+1}^{c_{m+2}},q_{m+1}^{c_{m+2}}+q_{m+1}^{c_{m+2}-1}]\cap\mathcal{P})\geq\frac{d_{3}q_{m+1}^{c_{m+2}-1}}{\log q_{m+1}^{c_{m+2}}},$
where $d_{3}$ is a small positive constant. In the case $m=k$, both (3.10) and (3.11) are true by (3.8) and (3.9). We assume that there exists a sequence $(q_{m})_{m=k}^{M}$ of prime numbers satisfying (3.10) and (3.11) for every $m=k,\ldots,M-1$, where $M\geq k+1$. Then by applying Lemma 3.4 with $X\coloneqq q_{M}^{c_{M+1}}$, $\eta\coloneqq(c_{M+1}-1)/c_{M+1}$, and $c\coloneqq c_{M+2}$, there exists a prime number $q_{M+1}$ satisfying (3.10) and (3.11) with $m=M$. By induction, we obtain an infinite sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying (3.10) and (3.11).
@@
for every $m\in\mathbb{N}$, and hence $\lim_{m\to\infty}q_{m}^{1/C_{m}}=w$ and $\lim_{m\to\infty}(q_{m}+1)^{1/C_{m}}=w^{\prime}$ exist. In addition, $w\in\mathcal{W}(c_{k})$ holds by (3.12). On the other hand, combining (3.7) and (3.8),
@@
(3.16) $p_{k}\leq\xi^{C_{k}}\leq(p_{k}^{c_{k+1}}+p_{k}^{\theta c_{k+1}})^{1/c_{k+1}}=p_{n}(1+p_{k}^{c_{k+1}(\theta-1)})^{1/c_{k+1}}=p_{k}+O(p_{k}^{c_{k+1}(\theta-1)+1}).$
@@
Remark 3.6.
We can observe the main difficulties of Question 1.4 in (3.16). If $c_{1}=c_{2}=\cdots=2$, then roughly speaking, (3.16) should be
$p_{k}\leq\xi^{2^{k}}<p_{k}+p_{k}^{2(1/2-1)+1}=p_{k}+1.$
It does not suffice to obtain (3.14). Further, by Matomäki’s construction, we do not know which specific prime number $q$ satisfies (3.4). This is also one of the main difficulties.
@@
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k\in\mathbb{N}}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.3. Let $\xi$ be the smallest element in $\mathcal{W}(c_{k})$. Then the following properties are true:
@@
Claim. We have $12\leq\ell\leq(b\theta_{b})^{-1}+$ 1.
@@
Case ($b\geq 5$): It follows that
@@
which is a contradiction to the claim. Therefore, $\xi$ is transcendental.
@@
Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. Suppose that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number. Let $\beta=\xi^{C_{m}}$ and $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ be the conjugates over $\mathbb{Q}$ of $\beta$. We set
@@
Let $\delta_{k}=\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}$ for every $k\geq m$. Let $k$ be a sufficiently large integer in $\mathcal{I}_{b}\cap[m,\infty)$. Since $t_{k}$ and $\beta_{1}^{C_{k}/C_{m}}$ are real numbers, $\delta_{k}$ is also a real number. Further, we recall that $|\beta_{j}|<1$ for every $2\leq j\leq\ell$, which implies that $\delta_{k}\in(-1/2,1/2)$ since $k$ is sufficiently large. If $-1/2<\delta_{k}\leq 0$, then $\xi^{C_{k}}=\beta_{1}^{C_{k}/C_{m}}=t_{k}-\delta_{k}$, and hence $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$. If $20<\delta_{k}<1/$ 2, then by (3.13), $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}$ holds for some constant $K_{0}>0$. Therefore,
@@
Since $t_{k}-p_{k}\in\mathbb{Z}$ and $-c_{k+1}\theta_{b}<0$, this is a contradiction by choosing $k$ large enough. ∎
@@
Let $b$, $(c_{k})_{k\in\mathbb{N}}$, and $\xi$ be as in Lemma 4.1. If $b\in\{3,4\}$, then there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$.
@@
By applying Lemma 4.2, for sufficiently large $k\in\mathcal{I}_{3}$, we have $p_{k}^{3}=p_{k+1}+3x_{1}x_{2}p_{k}$. By $x_{1}x_{2}\in\mathbb{Z}$, we see that $p_{k+1}$ is divisible by $p_{k}$. This is a contradiction since $p_{k}$ and $p_{k+1}$ are distinct prime numbers.
@@
Then $\beta_{2}^{d_{k}}\leq 0$. This is a contradiction since $\beta_{2}\in\mathbb{R}$ and $d_{k}$ is a positive even number. ∎
@@ 5. Further discussions
We will give several assumptions leading to the transcendency of $\xi_{3}$.
@@
Suppose that $\xi_{3}$ is algebraic. Then by Theorem 1.2, there exists $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. Let $p_{k}=\lfloor\xi_{3}^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$. By Fermat’s little theorem, $p_{k}^{3}\equiv p_{k}\mod 3$. Thus, by Remark 4.4, we have $p_{k+1}\equiv p_{k}\mod 3$ for every sufficiently large $k\geq m$. This is a contradiction. ∎
@@
which is a contradiction. ∎
Let $p_{n}$ be the $n$th prime number for every $n\in\mathbb{N}$. Cramér [5] conjectured that $p_{n+1}-p_{n}=O((\log p_{n})^{2})$. If this conjecture is true, we can take $\theta^{\prime}$ in Proposition 5.2 as an arbitrarily small positive constant. Therefore, assuming Cramér’s conjecture, $\xi_{3}$ is transcendental.

### v2

@@ 3. Lemmas and auxiliary results
Take any $k\in\mathbb{N}$. By the definition of $\xi$, we see that $p_{k}\leq\xi^{C_{k}}<p_{k}+1$, which implies that $p_{k}^{c_{k+1}}\leq\xi^{C_{k+1}}<(p_{k}+1)^{c_{k+1}}$. We recall that $c_{k+1}$ is a positive integer, and hence $p_{k}^{c_{k+1}}\leq\lfloor\xi^{C_{k+1}}\rfloor<(p_{k}+1)^{c_{k+1}}$. It is not true that $p_{k+1}=(p_{k}+1)^{c_{k+1}}-1$ since $p_{k+1}$ is a prime number and $c_{k+1}$ is an integer greater than or equal to $2$. Therefore, we obtain (3.3). ∎
@@
Let $\mathcal{I}_{b}$ be as in (3.2). There exists $k_{0}=k_{0}(\xi)>0$ such that for all $k\in\mathcal{I}_{b}\cap[k_{0},\infty)$, we have
@@
Before proving Lemma 3.6, we exhibit the following result given by Matomäki (see [15, Lemma 9] and [14]).
@@
There exist positive real numbers $d_{1}<1$ and $D$ such that, for every sufficiently large $x$ and every $\gamma\in[1/2,1]$, the interval $[x,2x]$ contains at most $Dx^{2/3-\gamma}$ disjoint intervals $[n,n+n^{\gamma}]$ for which
@@
By applying this theorem, Matomäki [15] constructed an element in $\mathcal{W}(2)$ as Theorem 3.2. It is truly astonishing because Mills’ construction suggests that we need the existence of a prime number in each interval $[n,n+n^{1/2}]$ to construct an element in $\mathcal{W}(2)$. This problem is strongly connected with Legendre’s conjecture which asserts that for every $n$ there exists a prime number between $n^{2}$ and $(n+1)^{2}$. It is unsolved and believed to be extremely difficult. However, Theorem 3.7 with $\gamma=1/2$ states that there exists a prime number $p\in[n,n+n^{1/2}]$ except for very few intervals $[n,n+n^{1/2}]$. Therefore, Matomäki discovered that we can construct a PRC if the exceptional intervals are very few. We apply Theorem 3.7 to prove the following lemma.
@@
Let $E$ and $c$ be real numbers with $2\leq c\leq E$. Let $d_{1}$ be as in Theorem 3.7. Let $\epsilon\in(0,1)$ and $d_{2}>0$. Then, there exists $X_{0}=X_{0}(E,\epsilon,d_{2})>0$ such that for all real numbers $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$, if we have
(3.5) #$\#([X,X+X^{\eta}]\cap\mathcal{P})\geq\frac{d_{2}X^{\eta}}{\log X},$
then there exists $q\in[X,X+X^{\eta}]\cap\mathcal{P}$ such that
@@
Let $D$ be as in Theorem 3.7. Let $X_{0}$ be a sufficiently large parameter depending only on $E$, $\epsilon$, and $d_{2}$. Take arbitrary real numbers $X\geq X_{0}$ and $\eta\in[1/2,1-\epsilon]$. We suppose that (3.5) is true. Then, let $R=[X,X+X^{\eta}]\cap\mathcal{P}$. Each interval $[q^{c},q^{c}+q^{c-1}]$ $(q\in R)$ is completely contained in $[X^{c},2X^{c}]$. Indeed, we have
@@
$\leq X^{c}(1+X^{\eta-1})^{E}(1+X^{-1})\leq 2X^{c}$
since $\eta\leq 1-\epsilon$, $X\geq X_{0}$, and $X_{0}$ is sufficiently large. Setting $\gamma=(c-1)/c$, we have $\gamma\in[1/2,1)$. By $\eta\geq 1/2$ and $c\geq 2$, the following inequalities are true:
$\frac{d_{2}X^{\eta}}{\log X}-DX^{c(2/3-\gamma)}$ $=\frac{d_{2}X^{\eta}}{\log X}-DX^{1-c/3}=\frac{X^{\eta}}{\log X}\left(d_{2}-DX^{1-\eta-c/3}\log X\right)$
(3.7) $\geq\frac{{X^{\eta}}}{\log X}\left(d_{2}-DX^{-1/6}\log X\right)>0,$
@@
Let $B$ be as in (3.1), let $\epsilon=\min(1/B,1/4)$, and let $d_{2}=\min(d_{0},d_{1})$, where $d_{0}$ and $d_{1}$ denote as in Theorems 2.1 and 3.7, respectively. Let $X_{0}=X_{0}(B,\epsilon,d_{2})$ be as in Lemma 3.8. We now take $k_{0}$ as a sufficiently large positive parameter depending only on $B$, $\epsilon$, and $d_{2}$. By the assumption (3.8), we find an integer $k$ in $\mathcal{I}_{b}\cap[k_{0},\infty)$ satisfying (3.8). By Lemma 3.5, the inequality $p_{k}^{c_{k+1}}>p_{k+1}$ does not hold, and hence we have
@@
(3.10) #$\#([q_{k}^{c_{k+1}},q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}]\cap\mathcal{P})\geq\frac{d_{0}q_{k}^{\theta c_{k+1}}}{\log q_{k}^{c_{k+1}}}.$
We may assume that $q_{k}^{c_{k+1}}\geq X_{0}=X_{0}(B,\epsilon,d_{2})$ since $k\geq k_{0}$ and $k_{0}=k_{0}(B,\epsilon,d_{2})$ is sufficiently large. By (3.10) and Lemma 3.8 with
$E\coloneqq B,\quad c\coloneqq c_{k+2}\in[2,B],\quad\epsilon\coloneqq\epsilon,\quad d_{2}\coloneqq d_{2},$
$X\coloneqq q_{k}^{c_{k+1}},\quad\eta\coloneqq\theta\in[1/2,3/4]\subseteq[1/2,1-\epsilon],$
there exists $q_{k+1}\in\mathcal{P}$ satisfying
(3.11) $q_{k}^{c_{k+1}}\leq q_{k+1}\leq q_{k}^{c_{k+1}}+q_{k}^{\theta c_{k+1}}(\leq q_{k}^{c_{k+1}}+q_{k}^{c_{k+1}-1}),$
(3.12) #$\#([q_{k+1}^{c_{k+2}},q_{k+1}^{c_{k+2}}+q_{k+1}^{c_{k+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q^{c_{k+2}-1}}{\log q^{c_{k+2}}},$
where the last inequality of (3.11) follows from $c_{k+1}=b\geq 3$ and $\theta=21/40$. Let us construct a sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying the following properties: for every $m\geq k$, we have
@@
(3.14) #$\#([q_{m+1}^{c_{m+2}},q_{m+1}^{c_{m+2}}+q_{m+1}^{c_{m+2}-1}]\cap\mathcal{P})\geq\frac{d_{1}q_{m+1}^{c_{m+2}-1}}{\log q_{m+1}^{c_{m+2}}}.$
In the case $m=k$, both (3.13) and (3.14) are true by (3.11) and (3.12). We assume that there exists a sequence $(q_{m})_{m=k}^{M}$ of prime numbers satisfying (3.13) and (3.14) for every $m=k,\ldots,M-1$, where $M\geq k+1$. Then by applying (3.14) with $m=M-1$ and Lemma 3.8 with
$E\coloneqq B,\quad c\coloneqq c_{M+2}\in[2,B],\quad\epsilon\coloneqq\epsilon,\quad d_{2}\coloneqq d_{2},$
$X\coloneqq q_{M}^{c_{M+1}},\quad\eta\coloneqq(c_{M+1}-1)/c_{M+1}\in[1/2,1-1/B]\subseteq[1/2,1-\epsilon],$
we find that there exists a prime number $q_{M+1}$ satisfying (3.13) and (3.14) with $m=M$. By induction, we obtain an infinite sequence $(q_{m})_{m=k}^{\infty}$ of prime numbers satisfying (3.13) and (3.14).
@@
for every $m\in\mathbb{N}$, and hence $\lim_{m\to\infty}q_{m}^{1/C_{m}}=\lim_{m\to\infty}(q_{m}+1)^{1/C_{m}}\eqqcolon w$ exist. In addition, $w\in\mathcal{W}(c_{k})$ holds by (3.15). On the other hand, combining (3.9) and (3.11), we have
@@
(3.19) $p_{k}\leq\xi^{C_{k}}\leq(p_{k}^{c_{k+1}}+2p_{k}^{\theta c_{k+1}})^{1/c_{k+1}}=p_{k}(1+2p_{k}^{c_{k+1}(\theta-1)})^{1/c_{k+1}}\leq p_{k}+2p_{k}^{c_{k+1}(\theta-1)+1}.$
@@
Let $b$ be an integer greater than or equal to $3$. Let $(c_{k})_{k=1}^{\infty}$ be a sequence of integers satisfying (1), (2), and (3) in Theorem 1.5. Let $\xi$ be the smallest element in $\mathcal{W}(c_{k})$. Then the following properties are true:
@@
Claim: We have $12\leq\ell\leq(b\theta_{b})^{-1}+$ 1.
@@
Case ($b\geq 5$): We now recall that $\theta_{b}=1-\theta-1/b=19/40-1/b$, and so
@@
a contradiction to the claim. Therefore, $\xi$ is transcendental.
@@
Let $b$, $(c_{k})_{k=1}^{\infty}$, and $\xi$ be as in Lemma 4.1. Suppose that there exists $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number. Let $\beta=\xi^{C_{m}}$ and $\beta_{1}=\beta,\beta_{2},\ldots,\beta_{\ell}$ be the conjugates over $\mathbb{Q}$ of $\beta$. We set
@@
Let $\delta_{k}=\sum_{j=2}^{\ell}\beta_{j}^{C_{k}/C_{m}}$ for every $k\geq m$. Let $k$ be a sufficiently large integer in $\mathcal{I}_{b}\cap[m,\infty)$. Since $t_{k}$ and $\beta_{1}^{C_{k}/C_{m}}$ are real numbers, $\delta_{k}$ is also a real number. Further, we recall that $|\beta_{j}|<1$ for every $2\leq j\leq\ell$, which implies that $\delta_{k}\in(-1/2,1/2)$ since $k$ is sufficiently large. If $20<\delta_{k}<1/$ 2, then by (3.16), $0\leq\xi^{C_{k}}-p_{k}\leq K_{0}p_{k}^{-c_{k+1}\theta_{b}}$ holds for some constant $K_{0}>0$. Therefore,
@@
Since we have $\delta_{k}>0$, $t_{k}-p_{k}\in\mathbb{Z}$, and $-c_{k+1}\theta_{b}<0$, this is a contradiction by choosing $k$ large enough. Thus, $-1/2<\delta_{k}\leq 0$ holds and $\xi^{C_{k}}=\beta_{1}^{C_{k}/C_{m}}=t_{k}-\delta_{k}$, and hence $t_{k}=\lfloor\xi^{C_{k}}\rfloor=p_{k}$. ∎
@@
Let $b$, $(c_{k})_{k=1}^{\infty}$, and $\xi$ be as in Lemma 4.1. If $b\in\{3,4\}$, then there does not exist $m\in\mathbb{N}$ such that $\xi^{C_{m}}$ is a Pisot number of degree $2$.
@@
By applying Lemma 4.2, for sufficiently large $k\in\mathcal{I}_{3}$, we have $p_{k}^{3}=p_{k+1}+3x_{1}x_{2}p_{k}$. By $x_{1}x_{2}\in\mathbb{Z}$, we see that $p_{k+1}$ is divisible by $p_{k}$, a contradiction since $p_{k}$ and $p_{k+1}$ are distinct prime numbers.
@@
Then $\beta_{2}^{d_{k}}\leq 0$, a contradiction since $\beta_{2}\in\mathbb{R}$ and $d_{k}$ is a positive even number. ∎
@@ 5. Further discussions
We give several assumptions leading to the transcendency of $\xi_{3}$.
@@
Suppose that $\xi_{3}$ is algebraic. Then by Theorem 1.2, there exists $m\in\mathbb{N}$ such that $\xi_{3}^{3^{m}}$ is a Pisot number of degree $3$. Let $p_{k}=\lfloor\xi_{3}^{3^{k}}\rfloor$ for every $k\in\mathbb{N}$. By Fermat’s little theorem, $p_{k}^{3}\equiv p_{k}\mod 3$. Thus, by Remark 4.4, we have $p_{k+1}\equiv p_{k}\mod 3$ for every sufficiently large $k\geq m$. ∎
@@
a contradiction. ∎
Let $p_{n}$ be the $n$th prime number for every $n\in\mathbb{N}$. Cramér [7] conjectured that $p_{n+1}-p_{n}=O((\log p_{n})^{2})$. If this conjecture is true, we can take $\theta^{\prime}$ in Proposition 5.2 as an arbitrarily small positive real number. Therefore, assuming Cramér’s conjecture, $\xi_{3}$ is transcendental.
Question 5.3.
Assuming the Riemann hypothesis, is Mills’ constant transcendental?
Carneiro, Milinovich, and, Soundararajan [5, Theorem 1.5] showed that for all real numbers $x\geq 4$, there exists a prime number $p\in\left(x,x+\frac{22}{25}\sqrt{x}\log x\right)$ if the Riemann hypothesis is true. However, this does not reach the assumption in Proposition 5.2. We need more arithmetic (or algebraic) properties of Mills primes to answer the question affirmatively.